OpenStudy (anonymous):

When two resistors are connected in parallel, the effective resistance is given by the expression 1/(1/R_1+1/R_2) Simplify the complex fraction

OpenStudy (kropot72):

The first step is to convert the denominator into a single fraction as follows: \[\large \frac{1}{R _{1}}+\frac{1}{R _{2}}=\frac{R _{1}+R _{2}}{R _{1}R _{2}}\] So the effective resistance is now given by: \[\large \frac{1}{\frac{R _{1}+R _{2}}{R _{1}R _{2}}}\ .........(1)\] The next step is to multiply the numerator and the denominator of (1) by the reciprocal of the denominator.

OpenStudy (kropot72):

The reciprocal of the denominator of (1) is: \[\large \frac{R _{1}R _{2}}{R _{1}+R _{2}}\]

OpenStudy (kropot72):

@TheKavMan Are you following the steps?

OpenStudy (anonymous):

@kropot72 so 1*(R1R2)/(R1+R2) correct?

OpenStudy (radar):

Product over sum. Is a simple method when calculating the total resistance of two resistors connected in parallel.

OpenStudy (anonymous):

Thanks alot :)

OpenStudy (kropot72):

The simplification is: \[\large R _{effective}=\frac{R _{1}R _{2}}{R _{1}+R _{2}}\]